# Sequential Change Detection with Differential Privacy

### v1

@@ Abstract
Index Terms:
Sequential change detection, differential privacy, CUSUM, average run length, detection delay.
@@ I Introduction
Sequential change detection is a fundamental problem in statistics and signal processing, with applications across a wide range of practical domains. The canonical formulation considers a sequence of observations sampled independently, with an unknown changepoint at which the underlying distribution switches from one distribution to an alternative. The goal of sequential change detection is to detect the occurrence of the distributional change with a minimal delay while controlling the false alarm rate. This problem is of major importance in many applications, such as seismic event detection [1], quality control [2], dynamical systems [3], healthcare [4], social networks [5], anomaly detection [6], detection of attacks [7], etc. However, in many of these applications, the data may contain sensitive personal information such as financial or medical records [8], while traditional sequential detection procedures typically release the decision directly without appropriate privacy protection. Thus, procedures with good detection ability while preserving individual’s information are highly desirable.
@@
Since Dwork et al.’s pioneering work [9], differential privacy has garnered much attention. A wide variety of differentially private procedures with theoretical efficiency guarantees have been developed for many statistical problems such as point estimation [8, 10] and hypothesis testing [11, 12, 13]. Informally, the field of differential privacy provides a systematic tool for constructing private algorithms or procedures by adding designed random noise, such that the output has a similar distribution whether or not the data is present for each individual participant, thereby helping to protect the information of individuals in the dataset.
@@
In this article, we study the problem of differentially private sequential change detection when both pre-change and post-change distributions are known and specified. We first extend the classical notion of $\epsilon$-differential privacy [9] for fixed databases to sequential change detection tasks over potentially infinite data streams, where the sample size is not fixed. We then develop a differentially private (DP) variant of the CUSUM procedure, called DP-CUSUM, which satisfies the new $\epsilon$-DP constraint. Our proposed DP-CUSUM procedure involves computing the CUSUM statistics while adding an independent Laplace noise at each time step, as well as to the detection threshold. By doing this, we do not release the CUSUM statistics or the threshold directly during the detection procedure. Meanwhile, the DP-CUSUM retains the same computational efficiency as the classical CUSUM procedure, as its computational overhead is minimal when noise is added at each time step.
@@
Our main contributions can be summarized as follows. First, we introduce a new concept of $\epsilon$-DP for sequential detection procedures when the sample size is not fixed. Second, we construct DP-CUSUM procedure that satisfies the new $\epsilon$-DP constraint. Third, under the assumption that the log-likelihood ratio statistic is always bounded by a known constant, we prove a lower bound to the average run length (ARL) of DP-CUSUM which enables an analytical way of selecting the detection threshold, and an upper bound to Lorden’s worst-case detection delay (WADD) [14], which thus proves its asymptotic optimality under weak privacy constraints. Fourth, we extend our results to scenarios with general unbounded log-likelihood ratios by relaxing to a slightly weaker differential-privacy definition with an added small constant $\delta$ in the DP definition. Finally, we validate our theoretical findings through numerical simulations, illustrating how the choice of privacy parameter affects the delay-versus-false-alarm trade-off in practical settings.
@@ I-A Related Work
The first optimality result for sequential change detection appears in [23], with a focus on detecting a change in the drift of a Brownian motion. It was studied under a Bayesian framework, where the changepoint was modeled as an independent, exponentially distributed random variable. In contrast, the non‑Bayesian (minimax) setting considers the changepoint to be deterministic but unknown. Under the minimax setting, the CUSUM procedure is arguably the most widely adopted change detection algorithm for the classical setup of detecting a change from a known distribution to a known alternative [15]. The CUSUM procedure was first proved to be asymptotically optimum in [14] when observations are i.i.d. before and after the change. Its exact optimality under the same data model was established in [24]. In addition to its strong optimality guarantees, the CUSUM statistic admits a recursive update, making it computationally efficient for sequential settings that require processing each new sample immediately. Although the problem of change detection when both pre-change and post-change distributions are known is fully studied in literature, these procedures do not consider the data privacy-preserving guarantees.
First introduced by Dwork et al. [9], differential privacy has garnered much attention. Many classical statistical procedures are tailored to satisfy the privacy guarantees [8, 10, 11, 12, 13]. In particular, in the area of sequential change detection, an $\epsilon$-DP procedure was proposed in [25] to estimate the change point for sequential data when both pre-change and post-change are fully specified. This method relies on a fixed‑size sliding window and applies an offline change‑point estimator within each window, which may lead to relatively higher memory and computational overhead. Extensions to unknown post-change distributions have been considered in the univariate setting in [26]. More recent efforts have explored broader data modalities and definitions of privacy. For instance, in [27], sequential change detection for multivariate nonparametric regression was studied under local differential privacy. In [28], the authors extended the analysis of privatized networks from static to dynamic, underscoring the complexities and challenges of preserving privacy in the dynamic analysis of network data. However, none of these work provide the theoretical detection performance of their proposed private procedures in terms of average run lengh and the worst-case detection delay, which are two fundamental but important properties for sequential change detection problem.
@@ II-A Basics of Classical Sequential Change Detection
In this article, we assume that $f_{0}$ and $f_{1}$ are both known. Here, $\tau\in\{0,1,2,\ldots\}$ is a deterministic but unknown change time and the goal is to raise the alarm as quickly as possible after the change has occurred, while properly controlling the false alarm rate [19, 17, 21].
A sequential change detection procedure consists of a stopping time $T$ which denotes the time we stop and declare that a change has occurred before time $T$. Here $T$ is an integer-valued random variable, and the decision $\{T=t\}$ is based only on the observations in the first $t$ time steps. That is, $\{T=t\}\subseteq\mathscr{F}_{t},$ where we define the filtration $\mathscr{F}_{t}=\sigma\{X_{1},\ldots,X_{t}\}$ and let $\mathscr{F}_{0}$ denote the trivial sigma-algebra. To evaluate the performance of $T,$ we further denote by $\mathbf{P}_{\infty}$ and $\mathbf{E}_{\infty}$ the probability measure and corresponding expectation when all samples follow the pre-change distribution $f_{0}$ (i.e., the change occurs at $\infty$). Similarly, we use $\mathbf{P}_{0}$ and $\mathbf{E}_{0}$ to represent the probability measure and expectation when all samples follow the post-change distribution $f_{1}$ (i.e., the change occurs at 0). More generally, we denote by $\mathbf{P}_{\tau}$ and $\mathbf{E}_{\tau}$ the probability measure and expectation when the change happens at time $\tau$. Under the classical minimax formulation for the sequential change detection problem [14, 24], the optimal detection procedure is the one solving the following constrained optimization problem:
@@
For completeness, we provide in the following Lemma an asymptotic expression for the performance of the CUSUM procedure, which also serves as the information-theoretic lower bound for the WADD in problem (2). The proof of the following Lemma can be found in [29, Lemma 1].
@@
$\mathbf{E}_{\infty}[T(b_{\gamma})]\geq\gamma,~~~~\mathbf{E}_{0}[T(b_{\gamma})]=\frac{\log\gamma}{I_{0}}(1+o(1)),$ (5)
@@
From Lemma 1 we conclude that by applying the CUSUM procedure defined in (4) with threshold $b=\log\gamma$, the corresponding CUSUM stopping time $T$ enjoys an asymptotic performance captured by (5). By the optimality of CUSUM procedure [24], no other stopping time $T^{\prime}$ that satisfies the same false alarm constraint can have a limiting value for the ratio $\mathbb{E}_{0}[T^{\prime}]/\frac{\log\gamma}{I_{0}}$ that is smaller than 1 as $\gamma\to\infty$.
@@ II-B Differentially Private Sequential Change Detection
While the classical CUSUM procedure achieves optimality in detecting distributional changes under the assumption of full data access, it may not be suitable in applications where individual-level data must remain private. In such scenarios, it is desirable to design change detection algorithms that also preserve the privacy of individuals contributing to the data stream. A principled way to achieve this is through the framework of differential privacy (DP) [9, 30].
@@
We should emphasize that in the definition of $\epsilon$-DP, the expectation is take over the randomness of the algorithm $\mathcal{A},$ while the two databases $D$ and $D^{\prime}$ are fixed. One common technique to achieve $\epsilon$-differential privacy is by adding a Laplace noise [30]. Specifically, for a real-valued deterministic function $L:\mathbb{R}^{n}\to\mathbb{R},$ define its sensitivity as $\Delta(L)=\max_{D,D^{\prime}\text{are neighbors}}|L(D)-L(D^{\prime})|.$ Then, a randomized $\epsilon$-DP algorithm can be obtained by adding an independent Laplace random noise $\text{Lap}(\Delta(L)/\epsilon)$ to the realization of the statistic $L(D).$ That is, $\tilde{L}(D)=L(D)+\text{Lap}(\Delta(L)/\epsilon)$ is a randomized algorithm satisfying the $\epsilon$-DP constraint. Here, we denote $\text{Lap}(\beta)$ as the Laplace random variable with zero mean and scale parameter $\beta>0$, with probability density function $f_{\text{Lap}(\beta)}(x)=\exp(-|x|/\beta)/(2\beta).$
In particular, for our problem of sequential change detection with positive integer output in $\mathbb{Z}^{+},$ we can modify the general definition of $\epsilon$-DP to obtain the following definition of $\epsilon$-DP sequential detection procedure. Let $X_{(1:t)}=(X_{1},\ldots,X_{t})$ denote the sequence of observations up to time $t$, and let $X^{\prime}_{(1:t)}$ be a neighboring dataset that differs from $X_{(1:t)}$ in exactly one entry (i.e., they differ in at most one $X_{i}$ for some $i\in[t]$). We now present the formal definition below.
@@
$\mathbf{P}_{\mathcal{T}}(\mathcal{T}=n\mid X_{(1:n)})\leq e^{\epsilon}\mathbf{P}_{\mathcal{T}}(\mathcal{T}=n\mid X^{\prime}_{(1:n)}),\quad\forall\,n\geq 1,$ (6)
@@
Here, a larger value of $\epsilon$ implies a weaker privacy constraint. Intuitively, the above definition means that altering any single observation in the data stream only slightly affects the distribution of the randomized stopping time $\mathcal{T}$. Therefore, one cannot easily infer individual data values from the detection output $\mathcal{T}$. Our goal is to devise a sequential detection procedure that satisfies the $\epsilon$-DP constraint in (6) while maintaining good detection performance with a small detection delay when the average run length is controlled. In the next section, we propose to first construct an $\epsilon$-DP sequential detection procedure based on the classical CUSUM procedure. Then, we will study the theoretical detection properties of our proposed procedure with the emphasis on how the privacy constraint parameter $\epsilon$ affects its detection performance.
@@ III Proposed DP-CUSUM Procedure with Bounded Log-Likelihood Ratio
$\Delta=\Delta(\ell):=\sup_{x,y\in\mathbb{R}^{k}}\left|\ell(x)-\ell(y)\right|=\sup_{x,y\in\mathbb{R}^{k}}|\log\frac{f_{1}(x)}{f_{0}(x)}-\log\frac{f_{1}(y)}{f_{0}(y)}|.$
@@
It is worthwhile mentioning that, compared to the existing private change detection procedure proposed in [25], we require a smaller Laplace noise $Z_{t},W\sim\text{Lap}(\frac{2\Delta}{\epsilon})$ added to both the CUSUM detection statistics and detection threshold, whereas the method in [25] added $Z_{t}\sim\text{Lap}(\frac{4\Delta}{\epsilon})$ to the detection statistics and $W\sim\text{Lap}(\frac{8\Delta}{\epsilon})$ to detection threshold. Such reduced noise can significantly improve the detection performance while still achieving $\epsilon$-DP, since smaller Laplace noise means less degradation in detection performance, as validated in our numerical comparisons in Section V.
Then, we analyze the theoretical detection performance of the proposed DP-CUSUM procedure $\tilde{T}(b)$. The following theorem provides a lower bound to the ARL of DP-CUSUM procedure.
@@
Assume $\Delta$ is bounded. For any $\epsilon>0$ and threshold $b>0$, define the following quantity
@@
The lower bound in (10) provides an analytical way to select the threshold without extensive simulation. That is, for any desired ARL $\gamma$, one can numerically solve for $b$ such that the lower bound in (10) is no smaller than $\gamma$. Furthermore, by Theorem 2, we can see by setting the threshold $b$ such that
$\frac{e^{h(\epsilon,\Delta)b-2}}{4(b+1)^{2}}=\gamma\Rightarrow b=b_{\gamma}=\frac{\log\gamma+O(\log\log\gamma)}{h(\epsilon,\Delta)}=\frac{\log\gamma}{h(\epsilon,\Delta)}(1+o(1)),$ (11)
@@
We are now ready to analyze an upper bound for the WADD of our proposed DP-CUSUM procedure. By Lemma 2, it suffices to establish an upper bound on $\mathbf{E}_{0}[\tilde{T}(b)]$. However, deriving such a bound is nontrivial due to the update rule, which includes both truncation at zero and the addition of Laplace noise at each time step. Since we only aim to find a suitable upper bound to the delay, we introduce an alternative stopping time in the proof that is guaranteed to incur a longer delay than $\tilde{T}(b)$ but is easier to analyze. The details can be found in Appendix B.
@@
Assume $\Delta$ is bounded. We have for any $b>8I_{0},$
@@
However, when $\epsilon<2\Delta$, the upper bound we established for the WADD of DP-CUSUM becomes $\frac{\log\gamma}{\epsilon/2\Delta\cdot I_{0}}(1+o(1))$, which is greater than the optimal detection delay $\frac{\log\gamma}{I_{0}}(1+o(1))$ of the classical (non-private) CUSUM. This implies that our proposed DP-CUSUM will still keep the data privacy under the stronger privacy constraint at the price of losing some detection efficiency. Moreover, it remains an open question what the tight information-theoretic lower bound for the WADD is under the $\epsilon$-DP constraint (6) when $\epsilon<2\Delta$, and whether DP-CUSUM retains asymptotic optimality in this regime. We leave this as a direction for future work.

### v2

@@ I Introduction
Sequential change detection is a fundamental problem in statistics and signal processing, with applications across a wide range of practical domains. The canonical formulation considers a sequence of observations sampled independently, with an unknown changepoint at which the underlying distribution switches from one distribution to an alternative. The goal of sequential change detection is to detect the occurrence of distributional change with minimal delay while controlling the false-alarm rate. This problem is of major importance in many applications, such as seismic event detection [29], quality control [22], dynamical systems [14], healthcare [2], social networks [16], anomaly detection [7], detection of attacks [25], etc. However, in many of these applications, the data may contain sensitive personal information such as financial or medical records [5], while traditional sequential detection procedures typically release the decision directly without appropriate privacy protection. Thus, procedures with good detection ability while preserving individuals’ information are highly desirable.
@@
Since Dwork et al.’s pioneering work [12], differential privacy has garnered much attention. A wide variety of differentially private procedures with theoretical efficiency guarantees have been developed for many statistical problems, such as point estimation [5, 1] and hypothesis testing [10, 6, 32]. Informally, differential privacy provides a systematic framework for constructing private algorithms or procedures by adding designed random noise, such that the output has a similar distribution regardless of whether the data are present for each individual participant, thereby helping to protect individuals’ information in the dataset.
@@
In this article, we study the problem of differentially private sequential change detection when both pre-change and post-change distributions are known and specified. We first extend the classical notion of $\epsilon$-differential privacy [12] from fixed databases to sequential change detection tasks over potentially infinite data streams, where the sample size is not fixed. We then develop a differentially private (DP) variant of the CUSUM procedure, called DP-CUSUM, which satisfies the new $\epsilon$-DP constraint. Our proposed DP-CUSUM procedure involves computing the CUSUM statistics while adding an independent Laplace noise at each time step, as well as to the detection threshold. By doing this, we do not release the CUSUM statistics or the threshold directly during the detection procedure. Meanwhile, the DP-CUSUM retains the same computational efficiency as the classical CUSUM procedure, as its computational overhead is minimal when noise is added at each time step.
@@
Our main contributions can be summarized as follows. First, we introduce a new concept of $\epsilon$-DP for sequential detection procedures when the sample size is not fixed. Second, we construct the DP-CUSUM procedure that satisfies the new $\epsilon$-DP constraint. Third, under the assumption that the log-likelihood ratio statistic is always bounded by a known constant, we prove a nonasymptotic lower bound to the average run length (ARL) of DP-CUSUM, which enables an analytical way of selecting the detection threshold, and a nonasymptotic upper bound to Lorden’s worst-case detection delay (WADD) [17]. These nonasymptotic results also imply the asymptotic optimality of our proposed DP-CUSUM procedure under weak privacy constraints. Fourth, we extend our results to scenarios with general unbounded log-likelihood ratios by relaxing to a slightly weaker differential-privacy definition that adds a small constant $\delta$ to the DP definition. Finally, we validate our theoretical findings through numerical simulations, illustrating how the choice of privacy parameter affects the delay-versus-false-alarm trade-off in practical settings.
@@ I-A Related Work
The first optimality result for sequential change detection appears in [23], which focuses on detecting a change in the drift of a Brownian motion. It was studied under a Bayesian framework, where the changepoint was modeled as an independent, exponentially distributed random variable. In contrast, the non‑Bayesian (minimax) setting considers the changepoint to be deterministic but unknown. Under the minimax setting, the CUSUM procedure is arguably the most widely adopted change detection algorithm for the classical setup of detecting a change from a known distribution to a known alternative [20]. The CUSUM procedure was first proved to be asymptotically optimum in [17] when observations are i.i.d. before and after the change. Its exact optimality under the same data model was established in [19]. In addition to its strong optimality guarantees, the CUSUM statistic admits a recursive update, making it computationally efficient for sequential settings that require processing each new sample immediately. Although the problem of change detection when both pre- and post-change distributions are known has been extensively studied in the literature, these procedures do not consider data privacy-preserving guarantees.
First introduced by Dwork et al. [12], differential privacy has garnered much attention. Many classical statistical procedures are tailored to satisfy the privacy guarantees [5, 1, 10, 6, 32]. In particular, in the area of sequential change detection, an $\epsilon$-DP procedure was proposed in [9] to estimate the change point for sequential data when both pre-change and post-change are fully specified. This method relies on a fixed-size sliding window and applies an offline change-point estimator within each window, which may result in higher memory and computational overhead. Extensions to unknown post-change distributions have been considered in the univariate setting in [8]. More recent efforts have explored broader data modalities and definitions of privacy. For instance, in [4], sequential change detection for multivariate nonparametric regression was studied under local differential privacy. In [15], the authors extended the analysis of privatized networks from static to dynamic, underscoring the complexities and challenges of preserving privacy in the dynamic analysis of network data. However, none of these work provide the theoretical detection performance of their proposed private procedures in terms of average run length and the worst-case detection delay, which are two fundamental but important properties for the sequential change detection problem.
@@ II-A Basics of Classical Sequential Change Detection
In this article, we assume that $f_{0}$ and $f_{1}$ are both known. Here, $\tau\in\{0,1,2,\ldots\}$ is a deterministic but unknown change time, and the goal is to raise the alarm as quickly as possible after the change has occurred, while properly controlling the false alarm rate [26, 21, 28].
A sequential change detection procedure consists of a stopping time $T$, denoting the time at which we stop and declare that a change has occurred before time $T$. Here $T$ is an integer-valued random variable, and the decision $\{T=t\}$ is based only on the observations in the first $t$ time steps. That is, $\{T=t\}\subseteq\mathscr{F}_{t},$ where we define the filtration $\mathscr{F}_{t}=\sigma\{X_{1},\ldots,X_{t}\}$ and let $\mathscr{F}_{0}$ denote the trivial sigma-algebra. To evaluate the performance of $T,$ we further denote by $\mathbf{P}_{\infty}$ and $\mathbf{E}_{\infty}$ the probability measure and corresponding expectation when all samples follow the pre-change distribution $f_{0}$ (i.e., the change occurs at $\infty$). Similarly, we use $\mathbf{P}_{0}$ and $\mathbf{E}_{0}$ to represent the probability measure and expectation when all samples follow the post-change distribution $f_{1}$ (i.e., the change occurs at 0). More generally, we denote by $\mathbf{P}_{\tau}$ and $\mathbf{E}_{\tau}$ the probability measure and expectation when the change happens at time $\tau$. Under the classical minimax formulation for the sequential change detection problem [17, 19], the optimal detection procedure is the one solving the following constrained optimization problem:
@@
For completeness, we provide in the following Lemma an asymptotic expression for the performance of the CUSUM procedure, which also serves as the information-theoretic lower bound for the WADD in problem (2). The proof of the following Lemma can be found in [31, Lemma 1].
@@
$\mathbf{E}_{\infty}[T(b_{\gamma})]\geq\gamma,~~~~\text{WADD}[T(b_{\gamma})]=\frac{\log\gamma}{I_{0}}(1+o(1)),$ (5)
@@
From Lemma 1 we conclude that by applying the CUSUM procedure defined in (4) with threshold $b=\log\gamma$, the corresponding CUSUM stopping time $T$ enjoys an asymptotic performance captured by (5). By the optimality of the CUSUM procedure [19], no other stopping time $T^{\prime}$ that satisfies the same false alarm constraint can have a limiting value for the ratio $\mathbb{E}_{0}[T^{\prime}]/\frac{\log\gamma}{I_{0}}$ that is smaller than 1 as $\gamma\to\infty$.
@@ II-B Differentially Private Sequential Change Detection
While the classical CUSUM procedure is optimal for detecting distributional changes under the assumption of full data access, it may be unsuitable for applications in which individual-level data must remain private. In such scenarios, it is desirable to design change detection algorithms that also preserve the privacy of individuals contributing to the data stream. A principled way to achieve this is through the framework of differential privacy (DP) [12, 13].
@@
We should emphasize that in the definition of $\epsilon$-DP, the expectation is taken over the randomness of the algorithm $\mathcal{A},$ while the two databases $D$ and $D^{\prime}$ are fixed. One common technique to achieve $\epsilon$-differential privacy is by adding a Laplace noise [13]. Specifically, for a real-valued deterministic function $L:\mathbb{R}^{n}\to\mathbb{R},$ define its sensitivity as $\Delta(L)=\max_{D,D^{\prime}\text{are neighbors}}|L(D)-L(D^{\prime})|.$ Then, a randomized $\epsilon$-DP algorithm can be obtained by adding an independent Laplace random noise $\text{Lap}(\Delta(L)/\epsilon)$ to the realization of the statistic $L(D).$ That is, $\tilde{L}(D)=L(D)+\text{Lap}(\Delta(L)/\epsilon)$ is a randomized algorithm satisfying the $\epsilon$-DP constraint. Here, we denote $\text{Lap}(\beta)$ as the Laplace random variable with zero mean and scale parameter $\beta>0$, with probability density function $f_{\text{Lap}(\beta)}(x)=\exp(-|x|/\beta)/(2\beta).$
In particular, for our problem of sequential change detection with positive integer output in $\mathbb{Z}^{+},$ we can modify the general definition of $\epsilon$-DP to obtain the following definition of $\epsilon$-DP sequential detection procedure. Let $X_{(1:t)}=(X_{1},\ldots,X_{t})$ denote the sequence of observations up to time $t$, and let $X^{\prime}_{(1:t)}$ be a neighboring dataset that differs from $X_{(1:t)}$ in exactly one entry (i.e., they differ in at most one $X_{i}$ for some $1\leq i\leq t$). We now present the formal definition below.
@@
$\mathbf{P}_{\mathcal{T}}(\mathcal{T}=n\mid X_{(1:n)})\leq e^{\epsilon}\mathbf{P}_{\mathcal{T}}(\mathcal{T}=n\mid X^{\prime}_{(1:n)}),\!~\forall\,n\geq 1,$ (6)
@@
Here, a larger value of $\epsilon$ implies a weaker privacy constraint. Intuitively, the above definition means that altering any single observation in the data stream only slightly affects the distribution of the randomized stopping time $\mathcal{T}$. Therefore, one cannot easily infer individual data values from the detection output $\mathcal{T}$. Our goal is to devise a sequential detection procedure that satisfies the $\epsilon$-DP constraint in (6) while maintaining good detection performance with a small detection delay when the average run length is controlled. In the next section, we propose to first construct an $\epsilon$-DP sequential detection procedure based on the classical CUSUM procedure. We then study the theoretical detection properties of our proposed procedure, with emphasis on how the privacy-constraint parameter $\epsilon$ affects its detection performance.
@@ III Proposed DP-CUSUM Procedure with Bounded Log-Likelihood Ratio
$\Delta=\Delta(\ell):=\sup_{x,y\in\mathbb{R}^{k}}\left|\ell(x)-\ell(y)\right|.$
@@
It is worthwhile mentioning that, compared to the existing private change detection procedure proposed in [9], we require a smaller Laplace noise $Z_{t},W\sim\text{Lap}(\frac{2\Delta}{\epsilon})$ added to both the CUSUM detection statistics and detection threshold, whereas the method in [9] added $Z_{t}\sim\text{Lap}(\frac{8\Delta}{\epsilon})$ to the detection statistics and $W\sim\text{Lap}(\frac{4\Delta}{\epsilon})$ to detection threshold. Such reduced noise can significantly improve the detection performance while still achieving $\epsilon$-DP, since smaller Laplace noise means less degradation in detection performance, as validated in our numerical comparisons in Section V. Intuitively, our proposed method achieves $\epsilon$-DP under a smaller noise scale because the DP-CUSUM procedure is analyzed as a single, unified mechanism without invoking privacy composition. We allocate the full privacy budget $\epsilon$ to the entire detection process and exploit the structural property of the recursive CUSUM statistic. In contrast, the method in [9] adopts a two-stage design that combines the AboveThreshold mechanism with the offline Report Noisy Max procedure. Its DP guarantee relies on the direct composition theorem, where each component consumes half of the total privacy budget ($\epsilon/2$). Consequently, to guarantee overall $\epsilon$-DP, the AboveThreshold component alone must use a larger noise scale corresponding to its allocated privacy budget of $\epsilon/2$.
Then, we analyze the theoretical detection performance of the proposed DP-CUSUM procedure $\tilde{T}(b)$. Note the randomized detection procedure $\tilde{T}(b)$ also depends on the added Laplace random variables $Z_{t}$ and $W.$ That is, the decision $\{\tilde{T}=t\}$ is based on the observations $\{X_{1},\cdots,X_{t}\}$ and the corresponding added Laplace noise $\{Z_{1},\cdots,Z_{t},W\}.$ Thus, throughout this paper, when analyzing the randomized detection procedure $\tilde{T}(b)$, we denote by $\mathbf{P}_{\infty}$ and $\mathbf{E}_{\infty}$ the joint probability measure and corresponding expectation when all samples $\{X_{t}\}_{t\geq 1}$ are i.i.d and follow the pre-change distribution $f_{0}$ (i.e., the change occurs at $\infty$), $\{Z_{t}\}_{t\geq 1}$ are i.i.d with Lap($\frac{2\Delta}{\epsilon}$), and $W$ follow Lap($\frac{2\Delta}{\epsilon}$). Similarly, we use $\mathbf{P}_{0}$ and $\mathbf{E}_{0}$ to represent the joint probability measure and expectation when all samples $\{X_{t}\}_{t\geq 1}$ follow the post-change distribution $f_{1}$ (i.e., the change occurs at 0) and $\{Z_{t}\}$ and W remain the same Laplace distribution. More generally, we denote by $\mathbf{P}_{\tau}$ and $\mathbf{E}_{\tau}$ the joint probability measure and expectation when the change happens at time $\tau$. Then, the ARL and WADD of our proposed procedure is defined by $\mathbf{E}_{\infty}[\tilde{T}(b)]$ and $\sup_{\tau\geq 0}\,\operatorname*{ess\,sup}\mathbf{E}_{\tau}[(\tilde{T}(b)-\tau)^{+}|X_{1},\cdots,X_{\tau}]$ respectively.
Then, the following theorem provides a lower bound to the ARL of the DP-CUSUM procedure.
@@
Assume $\Delta$ is bounded. For any $\epsilon>0$ and threshold $b>2,$ define the following quantity
@@
The lower bound in (10) provides an analytical method for selecting the threshold without extensive simulation. That is, for any desired ARL $\gamma$, one can numerically solve for $b$ such that the lower bound in (10) is no smaller than $\gamma$. Therefore, by Theorem 2, we can see by setting the threshold $b$ such that
$\frac{e^{h(\epsilon,\Delta)b-2}}{4(b+1)^{2}}=\gamma\Rightarrow b=b_{\gamma}=\frac{\log\gamma}{h(\epsilon,\Delta)}(1+o(1)),$ (11)
@@
We are now ready to analyze an upper bound for the WADD of our proposed DP-CUSUM procedure. By Lemma 2, it suffices to establish an upper bound on $\mathbf{E}_{0}[\tilde{T}(b)]$. However, deriving such a bound is nontrivial due to the update rule, which includes both truncation at zero and the addition of Laplace noise at each time step. Since we only aim to find a suitable upper bound to the delay, we introduce an alternative stopping time in the proof that is guaranteed to incur a longer delay than $\tilde{T}(b)$ but is easier to analyze. The details are provided in Appendix B.
@@
Assume $\Delta$ is bounded. We have for any $b>0$,
@@
However, when $\epsilon<2\Delta$, the upper bound we established for the WADD of DP-CUSUM becomes $\frac{\log\gamma}{I_{0}}(\frac{2\Delta}{\epsilon})(1+o(1))$, which is greater than the optimal detection delay $\frac{\log\gamma}{I_{0}}(1+o(1))$ of the classical (non-private) CUSUM. This implies that our proposed DP-CUSUM will still preserve data privacy under the stronger privacy constraint, albeit at the cost of some detection efficiency. This illustrates the fundamental trade-off between increased detection delay and privacy gain. Specifically, we can use the privacy parameter $\epsilon$ to quantify the privacy gain of a differentially private method. A lower $\epsilon$ value indicates a higher level of privacy, while a higher $\epsilon$ value provides less privacy but greater accuracy. The exact CUSUM procedure given in Eq. (4) can be viewed as the extreme case with $\epsilon=+\infty$ (i.e., no privacy at all), as the exact CUSUM procedure does not satisfy any finite differential privacy guarantee. The WADD in Eq. (14) establishes an explicit privacy–delay trade-off that links the delay directly to $\epsilon$, where the delay approaches that of the exact CUSUM as $\epsilon$ becomes sufficiently large. Moreover, it remains an open question what the tight information-theoretic lower bound for the WADD is under the $\epsilon$-DP constraint (6) when $\epsilon<2\Delta$, and whether DP-CUSUM retains asymptotic optimality in this regime. We leave this as a direction for future work.
Remark 1 (Relaxed DP for $n>n_{0}$).
The $\epsilon$-DP requirement in Eq. (6) may appear restrictive when the sample size $n$ is small (e.g., $n=1$). In some cases, privacy protection is required only for data collected after a certain point in time—for example, when a sequence begins with publicly available or non-sensitive warm-start data. In such settings, the DP condition in Eq. (6) can be relaxed to hold only for all $n>n_{0}$ for some positive integer $n_{0}$, thereby enforcing privacy guarantees only for newly arriving data. In such cases, our DP-CUSUM procedure in Algorithm 1 could be modified accordingly so that Laplace noise is added only when $n>n_{0}.$ That is, $\tilde{S}_{t}=S_{t}+Z_{t}\mathbbm{1}_{\{t>n_{0}\}}$. Note that in such cases, the proof techniques of Theorems 2 and 3 remain valid with only minor modifications that do not affect the asymptotic order of the ARL and WADD as $b\to\infty$, provided that $n_{0}$ is a fixed integer independent of $b$. Therefore, our asymptotic results and the argument for asymptotic optimality remain unchanged under this relaxed DP requirement.